OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {((-1, 11), (0, 8), (1, 7), (2, 5)} (-1, 11) (0, 8) (1, 7) (2, 5) @dusty

OpenStudy (anonymous):

this one is pretty tuff lol

OpenStudy (anonymous):

yeah I know right?

OpenStudy (anonymous):

really hard...

OpenStudy (raden):

trials and error :)

OpenStudy (anonymous):

what? lol

OpenStudy (raden):

it means u just trial one by one, if error miss it :P for example : for the point (-1,11) it is given x=-1 and y=11 so, just prove is that point on the line true y=-2x+9 11 = -2(-1) + 9 11 = 2 + 9 11 = 11 so, it is right

OpenStudy (anonymous):

what is the aswer, just so I can proove

OpenStudy (raden):

for (0, 8) ---> x=0, y = 8 y = -2x + 9 8 = -2(0) + 9 8 = 9 LHS not equal RHS this is incorrect :)

OpenStudy (raden):

now, can u prove for the point (1,7) ?

OpenStudy (anonymous):

7 = -2 + 9 7=7

OpenStudy (raden):

just prove is LHS equal or not equal with RHS

OpenStudy (raden):

so, that's right as the solution too

OpenStudy (raden):

the last point ?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks

OpenStudy (raden):

you're welcome